QUESTION 4

  1. If you take a 171 mL of a 1.4 M NaCl solution and dilute it to 500 mL, what is the molarity of the final solution?

    Enter the numerical answer in decimal notation

QUESTION 10

  1. Information - Colligative Properties

    Colligative properties of solutions are those properties which depend only on the number of dissolved solute particles in solution not the chemical properties of the solute. For example when a solute is dissolved in a solvent, vapor pressure depression of the solvent occurs, resulting in an increase its boiling point temperature (boiling point elevation) and a decrease in its freezing point temperature (freezing point depression). X420= 1... Vapor Pressure, torr Vapor Pressure, torr 0 5 25 30 10 15 20 Temperature, °C XH20 = 0.75 10 20 30 70 80 90 100 110

In the plot above, XH2O = 1 is the vapor pressure data for pure water and XH2O = 0.75 is the vapor pressure data for a solution of water and some non-volatile solute where Xsolute = 0.25. Why would depressing the vapor pressure result in increasing the boiling point of water?

The point at which the solution curve reaches 100oC, hence the boiling point, occurs at a higher temperature than does the curve for the pure solvent.

The point at which the solution curve reaches standard pressure, hence the boiling point, occurs at the same temperature as the curve for the pure solvent.

The point at which the solution curve reaches standard pressure, hence the boiling point, occurs at a higher temperature than does the curve for the pure solvent.

The point at which the solution curve reaches standard pressure, hence the boiling point, occurs at a lower temperature than does the curve for the pure solvent.

QUESTION 11

  1. Since a solute has the effect of lowering the freezing point as well as raising the boiling point, does the data in the graph suggest any linkage in the origin of the two effects?

    In the plot above, XH2O = 1 is the vapor pressure data for pure water and XH2O = 0.75 is the vapor pressure data for a solution of water and some non-volatile solute where Xsolute = 0.25.

    No, the graphs diverge from each other as the temperature increases.

    No, the plots are diverging as the vapor pressure increases.

    Yes, the two plots are converging near the origin.

    Yes, at 100oC, the vapor pressure for each plot is the same.

QUESTION 12

  1. Information - Mechanism of Vapor Pressure Depression

    The vapor pressure of a solution composed of a volatile solvent and nonvolatile solute is depressed due to the stabilization of the solution by the increased entropy (see Chapter 19) of the solution compared to the pure solvent. This lowers the vapor pressure at a given temperature

    If increasing the entropy of a solution stabilizes it, what does that suggest about the amount of heat input required to bring the solution to the condition of the pure solution?

    If the solution is stabilized, less heat input is required (ie the temperature has to be increased) to bring it back to the same condition as the pure solvent at lower temperature.

    If the solution is stabilized, more heat input is required (ie the temperature has to be decreased) to bring it back to the same condition as the pure solvent at lower temperature.

    If the solution is stabilized, less heat input is required (ie the temperature has to be decreased) to bring it back to the same condition as the pure solvent at lower temperature.

    If the solution is stabilized, more heat input is required (ie the temperature has to be increased) to bring it back to the same condition as the pure solvent at lower temperature.
Answer #1

4)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 1.4 M

V1 = 171 mL

V2 = 500 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (1.4*171)/500

M2 = 0.4788 M

Answer: 0.48 M

Only 1 question at a time please